# Uniform Feasibility For Smoothed Backup Control Barrier Functions

### v1

@@ Abstract
We study feasibility guarantees for safety filters developed using Control Barrier Functions (CBFs) when a safe set is defined using the pointwise minimum of continuously differentiable functions, a construction that is common for the backup CBF method and typically nonsmooth. We replace the minimum by its log–sum–exp (soft-min) smoothing and show that, under a strict safety condition, the smooth function becomes a CBF (or extended CBF) for a range of the smoothing parameter. For compact safe sets, we derive an explicit lower bound on the smoothing parameter that makes the smooth function a CBF and hence renders the corresponding safety filter feasible. For unbounded sets, we introduce tail conditions under which the smooth function satisfies an extended CBF condition uniformly. Finally, we apply these results to backup CBFs. We show that safety of a compact (terminal) backup set under a backup controller, together with a condition ensuring safety of the backup trajectories on the relevant boundary of the safe set, is sufficient for feasibility for backup CBFs. These results provide a recipe for a priori feasibility guarantees for smooth inner approximations of nonsmooth safe sets without the need for additional online certification.
@@ I Introduction
Safety is the main requirement for controlling modern autonomous systems. In safety-critical control, the objective is to provide a formal guarantee that, starting from a designated set, trajectories never leave this set during operation. Control Barrier Functions (CBFs) embed such guarantees as constraints in online optimization—often by penalizing the deviation from a desired control law, which leads to a design paradigm called safety filter [ames2019control]. The practicality of this framework ultimately hinges on the feasibility of the resulting optimization problem, which in turn requires a certificate that the constraint function is a valid CBF.
@@
Converse safety theorems study conditions such that a safe set admits a CBF, therefore they can be used to address the certification problem when the safe set is represented by a continuously differentiable function. Converse results mostly rely on compactness assumption of the set, and span stability-like conditions [prajna2005ontheNecessity], separation certificate via Meyer functions [wisniewski2016converse], and robust variants [ratschan2018converse]. Alternatives avoid compactness assumption at the expense of other restrictive conditions such as global Lipschitz constant [liu2022converse] or bounded Lie derivatives [gurriet2018towards]. Recently, a relaxed formulation of CBF, named extended CBF (eCBF), has been proposed to remove these assumptions from the converse side of necessity (of the existence of an eCBF), while retaining safety guarantees on the direct side of sufficiency [mestres2024converse].
@@
While converse safety theorems enable CBF certification, they presuppose the existence of a control law that keeps the system within the set, or equivalently, the set must be viable [^1] A set is viable if, from every point in it, the system can apply some admissible control that keeps the trajectory inside the set for all future time. under some control input in the sense of Aubin’s viability theory [aubin1991viability]. Ideally, one would find the largest viable set (viability kernel), but related computations are generally intractable in high dimensions due to the curse of dimensionality [mitchell2005aTime]. The backup control barrier function (BCBF) framework provides a scalable alternative [gurriet2018online, chen2021backup, wabersich2021predictive, breeden2022predictive]. Here, a backup controller is used to propagate states forward into a smaller viable “backup set,” and the set of all states that reach this backup set (at a terminal time) without leaving the safe region is implicitly defined via a minimum over time-indexed constraints, which leads to a nonsmooth. Since identifying a suitably small viable backup set can be achieved through convex relaxations or semidefinite programming, e.g., via LMIs [korda2013inner], this approach avoids direct computation of the viability kernel. The paradigm has been extended to data-driven settings via Koopman operators [folkestad2020data], to reduced-order models [molnar2023safety], to robust formulations [van2024disturbance], and to the multiple backup controllers setting [janwani2024learning]. However, feasibility guarantees in these BCBF-based methods often rely on ad hoc assumptions—such as presuming the existence of suitable class-$\mathcal{K}$ bounds without constructive certificates [gurriet2018online, chen2021backup], or imposing global compactness [gurriet2020scalable], or bounded Lie derivatives [rivera2024forward]—which limits systematic deployment.
@@
Contributions. We bridge BCBFs with converse safety certification to find feasibility guarantees for CBF-based safety filters. We do this by smoothing the nonsmooth pointwise minimum via the soft-min (log–sum–exp) operator and provide a range of parameters such that converse safety theorems in [mestres2024converse] can be employed for CBF/eCBF certification. In particular: (i) For compact safe sets, uniform continuity can be invoked to provide CBF conditions. (ii) For noncompact sets, we introduce explicit tail conditions under which the soft-min uniformly satisfies an eCBF condition. (iii) In the BCBF setting, we show that safety of a compact backup set under a backup controller, combined with a condition enforcing safety along the relevant boundary of the safe set, is sufficient to certify the nonsmooth backup CBF as a CBF. This relaxes the global compactness assumption in [gurriet2020scalable]. While the smoothing idea also appeared in [rabiee2025soft] for BCBFs, the approach of [rabiee2025soft] requires computing an inner feasible subset online via an optimization problem for given design parameters. In contrast, our results provide an a priori and uniform feasibility guarantee that eliminates the need for additional online certification.
@@
Organization. Section II reviews notation, (e)CBFs, converse safety, and the BCBF setup. Section III presents the soft-min construction and uniform thresholds for compact and noncompact cases. Section IV implements these results for BCBFs. Section V concludes the paper.
@@ II-B Control Barrier Functions and Converse Safety Theorems
Definition 2 (Extended CBF, [mestres2024converse]).
A function $h\in C^{1}$ is an extended CBF for (1) if there exists a function $\gamma\in\mathcal{K}_{\infty}\mathcal{K}$ with:
@@
CBFs and eCBFs imply that any locally Lipschitz continuous controller satisfying (3) or (4) ensures safety, provided that 0 is a regular value of $h$ [ames2019control, mestres2024converse]. One such controller that safeguards a desired (locally Lipschitz continuous) controller $k_{\rm des}$ is the safety filter:
$k_{\rm SF}(x)=$ $~\underset{u\in\mathbb{U}}{\rm argmin}~\|u-k_{\rm des}(x)\|$ (5)
$~\text{s.t.}~L_{f}h(x,u)\geq-\alpha(h(x)),$
where $\alpha(h(x))$ can be replaced with $\gamma(h(x),\|x\|)$ for eCBF. While certification of $h$ as a CBF or an eCBF implies the feasibility of the safety constraint, the optimization problem (5) becomes a QP (with efficient real-time solvers available) if $f$ is affine-in-control and $\mathbb{U}$ can be represented with affine constraints. Converse safety theorems certify $h$ as a CBF or an eCBF.
Theorem 1 (Converse safety for compact sets, [ames2019control]).
Let $\mathcal{S}$ be a compact set defined with $h\in C^{1}$ as in (2) and assume 0 is a regular value of $h$. If there exists a controller $k:\mathbb{R}^{n}\to\mathbb{U}$ such that $F(x)=f(x,k(x))$ is safe w.r.t. $\mathcal{S}$, then $h$ is a CBF.
Theorem 2 (Converse safety for general case, [mestres2024converse]).
Let $\mathcal{S}$ be a (potentially unbounded) set defined with $h\in C^{1}$ as in (2) and assume 0 is a regular value of $h$. If there exists a controller $k:\mathbb{R}^{n}\to\mathbb{U}$ such that the closed loop system $F$ is safe w.r.t. $\mathcal{S}$, then $h$ is an eCBF.
@@
The benefit of eCBFs with a more general class of functions $\gamma\in\mathcal{K}_{\infty}\mathcal{K}$ over $\alpha\in\mathcal{K}$ make them more suitable for unbounded sets. For example, even when $L_{F}h(x)$ becomes arbitrarily negative as $\|x\|\to\infty$, it is still plausible to construct a lower bound $\gamma(\cdot,\|x\|)$ using $L_{F}h$ thanks to the second term. However, missing this term, a CBF candidate may fail to admit a function $\alpha\in\mathcal{K}$. In particular, if $L_{F}h(x)\to$-$\infty$ as $\|x\|\to\infty$ and $h(x)=c\geq 0$ with a constant $c>0$, then no $\alpha\in\mathcal{K}$ can satisfy the barrier function condition with $F$ because $\alpha(c)$ is constant, see Example IV.1 in [mestres2024converse].
II-C Backup Control Barrier Functions
While converse safety theorems give us strong theoretical guarantees to certify any $h$ defining $\mathcal{S}$ as a CBF or an eCBF, they essentially require the existence of a control law rendering $\mathcal{S}$ safe. A scalable solution is proposed in [gurriet2018online]. This idea, which will be called backup control barrier functions in this paper, relies on constructing a (larger) safe set using finite-time safe trajectories that evolve within a (smaller) safe set $\mathcal{S}_{\rm b}\subseteq\mathcal{S}$.
@@ II-C Backup Control Barrier Functions
where
$h_{T}(x)\triangleq\min\left\{\min_{\tau\in[0,T]}h(\phi(x,\tau)),h_{\rm b}(\phi(x,T))\right\}.$ (8)
The set $\mathcal{S}_{T}$ contains all points $x\in\mathcal{S}$ such that, starting from $x$ at $\tau=0$, the system reaches $\mathcal{S}_{\rm b}$ by $\tau=T$ without leaving $\mathcal{S}$ at any $\tau\in[0,T]$ under the backup controller. Therefore, if $\mathcal{S}_{\rm b}$ is a safe set, so is $\mathcal{S}_{T}$, see [gurriet2020scalable, molnar2023safety] for a more formal statement.
The set $\mathcal{S}_{T}$ can potentially be used in converse safety theorems to certify $h_{T}$ as a CBF or an eCBF, which then guarantees the feasibility of the safety constraint in the safety filter with $h_{T}$. However, Theorems 1-2 require a continuously differentiable constraint function, whereas $h_{T}$ is not due to the minimum. To solve this problem, we will use a smooth inner approximation of $\min$ functions in $h_{T}$ in the next section. Before that, we use discretization for the minimum in $h_{T}$ over a continuous time interval $[0,T]$. In particular, consider a discretization $\tau_{i}\in\{0,\Delta\tau,\cdots,T\}$ with a uniform interval $\Delta\tau>0$ and $i\in\mathbb{I}_{N}$ with $N=T/\Delta\tau+1$. Discretization leads to an approximation of ${\mathcal{S}}_{T}$ defined with
@@
with an $\alpha_{i}\in\mathcal{K}$ for CBF and a $\gamma_{i}\in\mathcal{K}_{\infty}\mathcal{K}$ for eCBF. While (11) is sufficient for safety [^2] Technically, discretization yields a relaxation of the original safe set. While it is possible to shift the safety guarantee to $\hat{\mathcal{B}}$ by tightening constraints against imperfections emerging from discretization [gurriet2020scalable], we skip these details here to keep the focus on the merits of the approach presented. [he2025predictive], its feasibility depends on whether $\hat{b}$ can be certified as a CBF or an eCBF.
Our main goal in this paper is to obtain feasibility guarantees for CBF-based controllers such as (5) when the safety constraint is constructed as smoothened inner approximation of pointwise minimum of subconstraint functions. In the next section we set up this problem for a general setting and provide solutions for two separate cases: first, when the set $\hat{\mathcal{B}}$ is compact (using Theorem 1); then for a general case (using Theorem 2). Finally, we will investigate how these conditions translate back to the BCBF framework in Section IV.
@@ III-A Problem Formulation
When $h_{i}\in C^{1}$, MFCQ is satisfied whenever all the active gradients $\nabla h_{i}$ do not point in exactly opposite directions, so a common direction can increase all active constraints simultaneously.
@@
It is noted that the safety definition in Definition 3 is a stricter notion than the safety definition commonly used within the CBF framework [ames2019control]. In particular, inspired by the strict invariance in [aubin1991viability], we consider cases where any trajectory on the boundary cannot stay on the boundary, and it is strictly pushed inside. Defined on a stricter condition, Theorems 1-2 are still applicable for the safety definition in Definition 3.
According to the strict invariance result in Aubin [aubin1991viability, $\S 4.3$], $F(x)$ must belong to the relative interior of the contingent cone $T_{\hat{\mathcal{S}}}(x)$ for all $x\in\partial\hat{\mathcal{S}}$ to ensure strict safety. When $\hat{\mathcal{S}}$ is defined as in (12), and under MFCQ, the strict invariance condition becomes equivalent to [clarke1998nonsmooth]
@@
The strict positiveness of the Lie derivative $L_{F}\tilde{h}_{\theta}$ in (20) implies that $\nabla\tilde{h}_{\theta}(x)\!\neq\!0$ for all $x\in\partial\tilde{\mathcal{S}}_{\theta}$; so 0 is a regular value of $\tilde{h}_{\theta}$, and $\partial\tilde{\mathcal{S}}_{\theta}$ is a $C^{1}$ hypersurface. Therefore, (20) becomes sufficient for Aubin’s strict invariance condition [aubin1991viability], and Theorems 1-2 apply.
@@
While both our approach and that of [rabiee2025soft] employ the same type of smooth approximation (18) for the minimum operator (12), the nature of the resulting feasibility guarantees differs. In [rabiee2025soft], feasibility of the safety filter is only ensured on a subset $\Gamma\subseteq\tilde{\mathcal{S}}_{\theta}$, where $\Gamma$ is determined through an auxiliary optimization problem. In our work, we instead derive conditions under which feasibility holds on $\tilde{\mathcal{S}}_{\theta}$, and these conditions are obtained prior to implementation, without the need for additional online certification.

### v2

@@ Abstract
We study feasibility guarantees for safety filters developed using Control Barrier Functions (CBFs) when a safe set is defined using the pointwise minimum of continuously differentiable functions, a construction that is common for the backup CBF (BCBF) method and typically nonsmooth. We replace the minimum by its log–sum–exp (soft-min) smoothing and show that, under a strict safety condition, the smooth function becomes a CBF (or extended CBF) for a range of the smoothing parameter. For compact safe sets, we derive an explicit lower bound on the smoothing parameter that makes the smooth function a CBF and hence renders the corresponding safety constraint feasible. For unbounded sets, we introduce tail conditions under which the smooth function satisfies an extended CBF condition uniformly. Finally, we apply these results to BCBFs. We show that safety of a compact (terminal) backup set under a backup controller, together with a condition ensuring safety of the backup trajectories on the relevant boundary of the safe set, is sufficient for constraint feasibility for BCBFs. These results provide a recipe for a priori feasibility guarantees for smooth inner approximations of nonsmooth safe sets without the need for additional online certification.
@@ I Introduction
In safety-critical control, the objective is to provide a formal guarantee that, starting from a designated set, trajectories never leave this set during operation. Control Barrier Functions (CBFs) embed such guarantees as constraints in online optimization—often by penalizing the deviation from a desired control law, which leads to a design paradigm called safety filter [1]. The practicality of this framework ultimately hinges on the feasibility of the resulting optimization problem, which in turn requires a certificate that the constraint function is a valid CBF.
@@
Converse safety theorems study conditions such that a safe set admits a CBF, therefore they can be used to address the certification problem when the safe set is represented by a continuously differentiable function. Converse results mostly rely on compactness assumption of the set, and span stability-like conditions [17], separation certificate via Meyer functions [23], and robust variants [19]. Alternatives avoid compactness assumption at the expense of other restrictive conditions such as global Lipschitz constant [14] or bounded Lie derivatives [7]. Recently, a relaxed formulation of CBF, named extended CBF (eCBF), has been proposed to remove these assumptions from the converse side of necessity (of the existence of an eCBF), while retaining safety guarantees on the direct side of sufficiency [15].
@@
While converse safety theorems enable CBF certification, they presuppose the existence of a control law that keeps the system within the set. The backup control barrier function (BCBF) framework provides a scalable method to construct such a set [8, 4, 22, 3]. Here, a backup controller is used to propagate states forward into a smaller safe “backup (terminal) set,” and the set of all states that reach this backup set (at a terminal time) without leaving the safe region is implicitly defined via a minimum over time-indexed constraints, which leads to a nonsmooth. Identifying a suitably small backup sets can be achieved offline through Lyapunov equations, convex relaxations or semidefinite programming, e.g., via LMIs [12]. The paradigm has been extended to data-driven settings via Koopman operators [6], to reduced-order models [16], to robust formulations [21], and to the multiple backup controllers setting [11]. However, feasibility guarantees in these BCBF-based methods often rely on ad hoc assumptions—such as presuming the existence of suitable class-$\mathcal{K}$ bounds without constructive certificates [8, 4], or imposing global compactness [9], or bounded Lie derivatives [20]—which limits systematic deployment.
@@
Contributions. We bridge BCBFs with converse safety certification to find feasibility guarantees for CBF-based safety filters. We do this by smoothing the nonsmooth pointwise minimum via the soft-min (log–sum–exp) operator and provide a range of parameters such that converse safety theorems in [15] can be employed for CBF/eCBF certification. In particular: (i) For compact safe sets, uniform continuity can be invoked to provide CBF conditions, together with a formula to construct a class-$\mathcal{K}$ function guaranteeing the feasibility of the safety constraint. (ii) For noncompact sets, we introduce explicit tail conditions under which the soft-min uniformly satisfies an eCBF condition. (iii) In the BCBF setting, we show that safety of a compact backup set under a backup controller, combined with a condition enforcing safety along the relevant boundary of the safe set, is sufficient to certify a CBF. This relaxes the global compactness assumption in [9]. While the smoothing idea also appeared in [18] for BCBFs, the approach of [18] requires computing an inner feasible subset online via an optimization problem for given design parameters. In contrast, our result provides an a priori and uniform feasibility guarantee that eliminates the need for additional online certification. A numerical example illustrates the resulting design procedure end-to-end.
@@
Organization. Section II reviews notation, (e)CBFs, converse safety, and the BCBF setup. Section III presents the soft-min construction and uniform thresholds for compact and noncompact cases. Section IV implements these results for BCBFs. Section V demonstrates the approach on a numerical example. Section VI concludes the paper.
@@ II-B Control Barrier Functions and Converse Safety Theorems
Definition 2 (Extended CBF, [15]).
A function $h\in C^{1}$ is an extended CBF for (1) if there exists a function $\gamma\in\mathcal{K}_{\infty}\mathcal{K}$:
@@
CBFs and eCBFs imply that any locally Lipschitz continuous controller satisfying (3) or (4) ensures safety, provided that 0 is a regular value of $h$ [1, 15]. One such controller that safeguards a desired (locally Lipschitz continuous) controller $k_{\rm des}$ is the safety filter:
$k_{\mathrm{SF}}(x)=\underset{u\in\mathbb{U}}{\rm argmin}$ ${\color[rgb]{0,0,0}\|u-k_{\mathrm{des}}(x)\|_{2}^{2}}$ (5)
s.t. $L_{f}h(x,u)\geq-\alpha(h(x)),$
where $\alpha(h(x))$ can be replaced with $\gamma(h(x),\|x\|)$ for eCBF, and $\|\cdot\|_{2}$ denotes the Euclidean norm. While certification of $h$ as a CBF or an eCBF implies the feasibility of the safety constraint, the optimization problem (5) becomes a QP (with efficient real-time solvers available) if $f$ is affine-in-control and $\mathbb{U}$ can be represented with affine constraints. Converse safety theorems certify $h$ as a CBF or an eCBF.
Theorem 1 (Converse safety for compact sets, [1]).
Let $\mathcal{S}$ be a compact set defined with $h\in C^{1}$ and let 0 be a regular value of $h$. If there exists a controller $k:\mathbb{R}^{n}\to\mathbb{U}$ such that $F(x)=f(x,k(x))$ is safe w.r.t. $\mathcal{S}$, then $h$ is a CBF.
Theorem 2 (Converse safety for general case, [15]).
Let $\mathcal{S}$ be a (potentially unbounded) set defined with $h\in C^{1}$ and let 0 be a regular value of $h$. If there exists a controller $k:\mathbb{R}^{n}\to\mathbb{U}$ such that the closed loop system $F$ is safe w.r.t. $\mathcal{S}$, then $h$ is an eCBF.
@@
The benefit of eCBFs with a more general class of functions $\gamma\in\mathcal{K}_{\infty}\mathcal{K}$ over $\alpha\in\mathcal{K}$ make them more suitable for unbounded sets. For example, even when $L_{F}h(x)$ becomes arbitrarily negative as $\|x\|\to\infty$, it is still plausible to construct a lower bound $\gamma(\cdot,\|x\|)$ using $L_{F}h$ thanks to the second term. However, missing this term, a CBF candidate may fail to admit a function $\alpha\in\mathcal{K}$, see Example IV.1 in [15].
II-C Backup Control Barrier Functions (BCBF)
While converse safety theorems give us strong theoretical guarantees to certify any $h$ defining $\mathcal{S}$ as a CBF or an eCBF, they essentially require the existence of a control law rendering $\mathcal{S}$ safe. A solution is proposed in [8]. This idea, which will be called backup control barrier functions (BCBF) in this paper, relies on constructing a (larger) safe set using finite-time safe trajectories that evolve within a (smaller) safe set $\mathcal{S}_{\rm b}\subseteq\mathcal{S}$.
@@ II-C Backup Control Barrier Functions (BCBF)
where $h_{T}(x)\triangleq\min\left\{\min_{\tau\in[0,T]}h(\phi(x,\tau)),h_{\rm b}(\phi(x,T))\right\}.$ The set $\mathcal{S}_{T}$ contains all points $x\in\mathcal{S}$ such that, starting from $x$ at $\tau=0$, the system reaches $\mathcal{S}_{\rm b}$ by $\tau=T$ without leaving $\mathcal{S}$ at any $\tau\in[0,T]$ under the backup controller. Therefore, if $\mathcal{S}_{\rm b}$ is a safe set, so is $\mathcal{S}_{T}$, see [9, 16] for a more formal statement.
The set $\mathcal{S}_{T}$ can potentially be used in converse safety theorems to certify $h_{T}$ as a CBF or an eCBF, which then guarantees the feasibility of the safety constraint with $h_{T}$. However, Theorems 1-2 require a continuously differentiable constraint function, whereas $h_{T}$ is not due to the minimum. To solve this problem, we will use a smooth inner approximation of $\min$ functions in $h_{T}$ in the next section. Before that, we use discretization for the minimum in $h_{T}$ over a continuous time interval $[0,T]$. In particular, consider a discretization $\tau_{i}\in\{0,\Delta\tau,\cdots,T\}$ with a uniform interval $\Delta\tau>0$ and $i\in\mathbb{I}_{N}$ with $N=T/\Delta\tau+1$. Discretization leads to an approximation of ${\mathcal{S}}_{T}$ defined with
@@
with $\alpha_{i}\in\mathcal{K}$ for the CBF case and a $\gamma_{i}\in\mathcal{K}_{\infty}\mathcal{K}$ for the eCBF case. While constraints (10) are sufficient for safety [^1] Technically, discretization yields a relaxation of the original safe set. While it is possible to shift the safety guarantee to $\hat{\mathcal{B}}$ by tightening constraints against imperfections emerging from discretization [9], we skip these details here to keep the focus on the merits of the approach presented. [10], their feasibility depends on whether $\hat{b}$ can be certified as a CBF or an eCBF.
Our main goal in this paper is to obtain feasibility guarantees for the safety constraint of CBF-based controllers such as (5) when the constraint is constructed as smoothed inner approximation of pointwise minimum of subconstraint functions. In the next section we set up this problem for a general setting and provide solutions for two separate cases: first, when the set $\hat{\mathcal{B}}$ is compact (using Theorem 1); then for a general case (using Theorem 2). Finally, we will investigate how these conditions translate back to the BCBF framework in Section IV.
@@ III-A Problem Formulation
When $h_{i}\in C^{1},~\forall i\in\mathbb{I}_{N}$, MFCQ is satisfied whenever all the active gradients $\nabla h_{i}$ do not point in exactly opposite directions, so a common direction can increase all active constraints simultaneously.
@@
It is noted that the safety definition in Definition 3 is a stricter notion than the safety definition commonly used within the CBF framework [1]. In particular, inspired by the strict invariance in [2], we consider cases where any trajectory on the boundary cannot stay on the boundary, and it is strictly pushed inside. Defined on a stricter condition, Theorems 1-2 are still applicable for the safety definition in Definition 3.
According to the strict invariance result in Aubin [2, $\S 4.3$], $F(x)$ must belong to the relative interior of the contingent cone $T_{\hat{\mathcal{S}}}(x)$ for all $x\in\partial\hat{\mathcal{S}}$ to ensure strict safety. When $\hat{\mathcal{S}}$ is defined as in (11), and under MFCQ, the strict invariance condition becomes equivalent to [5]
@@
The strict positiveness of the Lie derivative $L_{F}\tilde{h}_{\theta}$ in (19) implies that $\nabla\tilde{h}_{\theta}(x)\!\neq\!0$ for all $x\in\partial\tilde{\mathcal{S}}_{\theta}$; so 0 is a regular value of $\tilde{h}_{\theta}$, and $\partial\tilde{\mathcal{S}}_{\theta}$ is a $C^{1}$ hypersurface. Therefore, (19) becomes sufficient for Aubin’s strict invariance condition [2], and Theorems 1-2 apply.
@@
The soft-min smoothing yields an inner approximation $\tilde{\mathcal{S}}_{\theta}\subseteq\hat{\mathcal{S}}$, and the gap can be quantified by the two-sided bound $\hat{h}(x)-\frac{\log N}{\theta}\leq\tilde{h}_{\theta}(x)\leq\hat{h}(x),\quad\forall x\in\hat{\mathcal{S}}$. So, the loss in the certified safe set vanishes as $\theta\to\infty$. While this conservativeness is the price paid for replacing the nonsmooth $\hat{h}$ with a $C^{1}$ function, it is precisely this smoothness that enables the feasibility guarantees developed in the next section.
Remark 3.
While both our approach and that of [18] employ the same type of smooth approximation (17) for the minimum operator (11), the nature of the resulting feasibility guarantees differs. In [18], feasibility of the safety filter is only ensured on a subset $\Gamma\subseteq\tilde{\mathcal{S}}_{\theta}$, where $\Gamma$ is determined through an auxiliary optimization problem. In our work, we instead derive conditions under which feasibility holds on $\tilde{\mathcal{S}}_{\theta}$, and these conditions are obtained prior to implementation, without the need for additional online certification.